• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name moussiegt

    Log entry time 08:01:04 on June28,2002

    Entry number 86935

    keyword=shift summary

    Last shift, with a lot of things to do...

    12h00 : end of production run 1861 (last of Q2_l50 kin) with 560 kevents.
    12h00 : start efficiency run 1862 (1 Mevents).
    12h10 : start changing kinematic to go to i40 by moving right arm only
    and changing momentum. Move OK but we had a interlock and quench fault in Q2 (right arm), whose current goes back to 0 A. So we paged tech on call.
    12h30 : We phoned him and he said we need to go to the hall to manually reboot it. So we asked for controled access. I will go to the hall and will phone him again to reboot Q2Right.
    01h05 : Q2 reseted. Current is ramping to 1400 A and we will set it to the right value.
    01h15 : Asked for beam back at 100 muA.
    ROC 14 crash when starting new run, so we had to reboot coda and
    ROC 14.
    01h30 : Start normalization run 1864.
    01h40 : Changing settings to Q2_i40 (left atm position and momentum).
    02h00 : Optics run started. (run 1865).
    02h10 : Start production run 1866. (kin i40).
    03h05 : Q1 (left arm) had He level fault and tripped. Current is going down to 0A.
    In the same time, the current decrease in the left dipole, seems because of low He level too. It says M3/M4 over temp fault (for the dipole).
    Run 1866 stopped at 850 kevents.
    So we reseted Q1. Dipole seemed to be trying to go back to the good value.
    We decided to let it and watch before paging tech on call.
    03h25 : Q1 is ready and so is the dipole, so we start a new run (1867).
    ROC 14 problems, reset it.
    03h30 : everything OK, start run 1868.
    03h45 : ROC 3 crash, so we decided to kill coda and reboot everything to prevent new problems for some time. Run 1868 with 290 kevents.
    03h55 : coda ready after we rebooted everything.
    Run 1869 started.
    04h20 : Noticed that coda is dead (but all components seems to work good).
    So killed it - run 1869 has 200 kevents.
    04h25 : start of run 1870.
    04h55 : stopped run 1870 with 500 kevents to change kinematic.
    05h00 : start efficiency run 1872.
    05h05 : start moving right spectrometer.
    05h10 : Q1 He Level fault.
    05h55 : all magnets Ok. start normalisation run 1873.
    06h15 : stopped run 1874 (1873 was junk) with 900 kevents.
    start changing kinematic.
    06h30 : seemed that Q1 left arm tripped during momentum change, so we will reset it.
    06h55 : magnets OK, start optics run 1875.
    07h00 : move to LD2 target for coincidence data.
    07h05 : start of run 1876 (production).
    07h25 : end of run 1876 (1 Mevents). start efficiency run 1877.
    07h30 : moving to LH2 target.
    07h45 : coincidence with hydrogen starts.
    07h45 : we noticed there was no count in T3 and T4 and then discovered all HV in left arm detectors were off. So we put them back on.
    08h00 : end of run 1878 (hydrogen run).

    end of shift.
    end of experiment.